# BigGait: Learning Gait Representation You Want by Large Vision Models

### 29 Feb 2024 v1

@@ 3.4 Implementation Details
$L=L_{tri}+L_{ce}+\gamma_{rec}L_{rec}+\gamma_{smo}L_{smo}+\gamma_{div}L_{div},$ (5)
where the first two terms are recognition losses of the downstream gait model, $L_{rec}$ is the mask reconstruction of the mask branch, and the last two are the smoothness and diversity losses of the denoising branch. In our implementations, $\gamma_{rec},\gamma_{smo}$ and $\gamma_{div}$ is respectively set as $1.0$, $0.01$, and $5.0$.
@@
Table 3: Rank-1 accuracy on CCPG: BigGait v.s. recent SoTA models using different inputs. $\text{GaitBase}^{p}$: human parsing inputs. $\text{GaitBase}^{p+s}$: parsing and silhouette inputs. BigGait30 represent that using 30 frames per seq to train BigGait instead of 8 frames.
@@
BigGait Ours 76.0 79.1 84.2 93.0 83.1
BigGait-L 79.0 82.3 86.7 94.5 85.6
[…]
Training. All images are processed with Pad-and-Resize to the size of $448\times 224$. We use the SGD optimizer with a weight decay of $0.0005$, a momentum of $0.9$ and an initial learning rate of $0.1$. The learning rate is reduced by a factor of $0.1$ at the $15$k, $25$k, $30$k and $35$k iterations, and the total number of iterations is $40$k. The batch size $(p,k,l)$ is set to $(8,8,8)$, where $p$ is the number of IDs, $k$ the number of sequences per ID, and $l$ the number of frames per sequence.
Table 4: Rank-1 accuracy on three authoritative datasets: BigGait v.s. recent SoTA methods. These models’ cross-domain performance, independently trained on CCPG (left), CASIA-B* (middle), and SUSTech1K (right).
(a). Trained on CCPG
@@
GaitSet [3] 47.4 40.9 25.8 8.2 11.0 12.8
GaitBase [13] 59.1 52.7 30.4 9.5 13.1 16.8
AP3D [18] 53.7 46.2 11.9 36.2 51.6 55.3
PSTA [45] 49.7 42.3 8.8 25.7 29.4 40.6
BigGait 84.1 77.5 33.8 46.0 59.0 64.7
(b). Trained on CASIA-B*
@@
GaitSet 10.6 16.4 17.2 24.9 7.2 12.2 12.8
GaitBase 10.6 18.1 21.4 28.7 8.1 11.8 15.6
AP3D 2.1 2.9 3.9 6.1 29.3 47.4 48.3
PSTA 1.7 1.9 3.4 5.0 19.9 37.5 34.6
BigGait 6.0 18.4 12.7 44.5 42.7 60.1 66.1
(c). Trained on SUSTech1K
@@
GaitSet 14.0 23.7 20.3 43.2 63.3 50.8 26.4
GaitBase 16.8 21.7 26.0 42.7 73.1 61.2 28.2
AP3D 5.5 7.9 13.9 35.1 56.7 48.1 15.3
PSTA 3.7 5.7 9.5 26.5 31.2 27.7 10.6
BigGait 6. 13.7 15.3 54.9 95.4 90.6 21.5
@@ 4.2 Main Results
This section presents a comprehensive study to show the effectiveness and generalization of BigGait on various datasets. For more efficient training, BigGait stands for BigGait-S as our baseline.
@@
Self-domain Evaluation. We compare BigGait with various SoTA methods, such as the skeleton-based [43, 48, 35], silhouette-based [3, 10, 25, 13], and end-to-end gait recognition methods [28], and video-based ReID methods [18, 45, 47], on the challenging CCPG. Besides, BigGait performs well on CASIA-B* and SUSTech1K, and these self-domain results are provided in the Supplementary Material.
@@
As shown in Tab. 3, BigGait achieves significantly better performance than the other SoTA methods on CCPG. We measure BigGait’s effectiveness in filtering out gait-irrelevant noises from two aspects.
@@
First, compared to video-based ReID methods [18, 45, 47], BigGait outperforms them considerably, e.g., +22.6% for full-changing (CL), +21.8% for ups-changing (UP), +14.5% for pants-changing (DN) and +1.6% for bag-changing (BG) scenarios. Accordingly, we consider that BigGait can efficiently extract robust identity representations from raw RGB videos. Fig. 5 will provide intuitive evidence by visualizing the activation difference between BigGait and video-based ReID methods.
@@
Second, compared to silhouette-based gait methods [3, 10, 25, 13], which perform well due to immunity to color/texture noises, BigGait still outperforms them by +4.4%, +4.1%, +7.4%, and +14.4% in CL, UP, DN, and BG settings. This indicates that BigGait’s immunity to gait-irrelevant noises is comparable to silhouette-based ones.
@@
Based on these analyses, we consider that gait features are the primary characteristics extracted by BigGait, even if gait-unrelated noises cannot be absolutely eliminated from RGB inputs. We further validate this point through cross-domain and visualization experiments.
@@
Cross-domain Evaluation. Tab. 4 shows cross-domain performance comparisons. We observe that BigGait’s cross-domain performance varies depending on the training set used. When trained on CCPG, BigGait exhibits strong adaptability to unseen datasets, i.e., outperforming both video-based ReID methods [18, 45] and silhouette-based methods [13, 3]. When the training set is CASIA-B*, BigGait displays less impressive performances in some settings, i.e., performing well in the ups-changing (UP), but poorly in the full-changing (CL) and pants-changing (DN). When crossing domains from SUSTech1K, BigGait struggles in more cross-dressing settings (CL, UP and DN).
@@
We consider that BigGait’s power in filtering gait-unrelated noises is impacted by the training data biases. According to sampling statistics, SUSTech1K only has 4.7% cross-dressing sequence pairs, causing BigGait to incorrectly learn unchanged clothing elements rather than subtle gait patterns. Fortunately, with the enrichment of training data, BigGait can self-correct this mistake. CASIA-B* includes 35.5% cross-dressing sample pairs, all of which are ups-changing, which improve BigGait’s resistance to ups-changing (UP), as shown in Tab. 4 (b). CCPG has more diverse outfits than others, containing 96.5% cross-dressing pairs, with 78.5% for ups-changing, 80.0% for pants-changing, and 62.0% for full-changing. As shown in Tab. 4 (a), BigGait can learn robust gait representations by training on CCPG. These analyses indicate that the distribution of training data can significantly influence BigGait’s outcomes, i.e, the more cross-dressing changes given, the more cross-dressing capacity obtained.
Figure 5: Visualization of activation maps. The left three are video-based ReID methods used for reference, and the left six are trained on CCPG.
@@
Visualization of Activation Maps. To delve deeper into the gait representation learned by BigGait, Fig. 5 shows visualizations of activation maps from both BigGait and video-based ReID methods [18, 45, 47], driven by the popular Grad-CAM [37] algorithm. These maps are from the first layer in BigGait’s downstream model, second layer in AP3D [18] and PSTA [45], and fourth block in PiT [47]. Four key insights emerge: a) Unlike video-based ReID methods distracted by cluttered backgrounds, BigGait solely attends to the body. b) The video-based ReID methods are vulnerable to high-frequency clothing texture, as shown in the red box of Fig. 5. c) $f_{ap}$ and $f_{de}$ target different body regions, with $f_{ap}$ emphasizing limbs and head via parsing-like priors, and $f_{de}$ highlighting human contours since suppressing texture noises on body. Fusing $f_{ap}$ and $f_{de}$ using attention weights to obtain $f_{fusion}$. d) As mentioned before, CASIA-B* only contains ups-changes, and SUSTech1K has few cloth-changes. Thus BigGait incorrectly highlights unchanged pants and clothing when trained on CAISA-B* and SUSTech1K, respectively. More visualizations are provided in Supplementary Material.
@@
4.3 Abalation Study
Impact of Pad-and-Resize. Compared (a) and (b) in Tab. 5, the Pad-and-Resize strategy is helpful for BigGait by preserving authentic body proportions.
Ablation of Each Branch. As shown in Tab. 5 (c), without the GRE module, the performance of BigGait drops to be similar to the video-based ReID method [18] in Tab. 3. This highlights the effectiveness of GRE module. Compared (a) to (d-f) in Tab. 5, each branch contributes substantially. This implies a mutually beneficial relationship among them. As shown in Tab. 5 (g), BigGait’s foreground mask is less discriminable for recognition compared to the silhouette in Tab. 3. Hence, we consider that BigGait’s foreground masks is approximate regions rather than precise body segmentations. This indicates that cutting-edge results in Tab. 5 (a) are achieved even without finely segmented foregrounds.
Ablation of Denoising Branch. Comparing (a) to (b-d) in Tab. 6, smoothness loss alone leads to a trivial solution, while combining it with diversity constraint improves results. Compared (a) with (e) and (f) in Tab. 6, we explore an appropriate channel number for the denoising branch.
Ablation of Upstream and Downstream Models. Compared (c) with (d) in Tab. 7, we verify that the gains primarily come from the knowledge embedded in DINOv2 rather than its model architecture. Compared (c) and (e) in Tab. 7, the performance suffers when fine-tuning the upstream DINOv2. Further, we explore the influence of different upstream/downstream models for BigGait. Compared (c) and (f) in Tab. 7, although the performance drops when changing BigGait’s downstream model, it still surpasses its silhouette-based counterpart in (a). As shown in Tab. 7 (g), BigGait remains effective even when changing BigGait’s upstream model, i.e., replacing DINOv2 [34] with SAM [23]. In most cases, it still outperforms its silhouette-based counterpart in (b). This highlights BigGait’s flexibility for upstream/downstream models.
@@ 4.3 Abalation Study
(d) DINOv2Without ${}^{\text{Without Pretraining}}$ GaitBase 36.6 44.0 62.4 79.3
@@
5 Challenges and Limitations
Challenges for LVMs-based Gait Recognition. Beyond attractive profits brought by BigGait, this paper also reveals two primary challenges for LVMs-based gait recognition: a) Interpretability. While we have introduced several explicit human gait priors as soft constraints, the learned gait representations, in comparison to conventional ones defined by clear and intuitive physical attributes, remain partially understandable. b) Purity. In gait methods that directly utilize RGB videos as input, a recurring challenge involves effectively reducing gait-unrelated noise within walking sequences. This task becomes even more demanding when attempting to preserve the purity of gait characteristics in LVMs-based gait recognition without explicit supervision.
Limitations of this paper. a) The influence of different upstream LVMs for BigGait is only preliminarily explored. This issue is worth a continuous study. b) The GRE module lacks of spatial-temporal designs.

### 22 Mar 2024 v2

@@ 3.4 Implementation Details
$L=L_{tri}+L_{ce}+\gamma_{rec}L_{rec}+\gamma_{smo}L_{smo}+\gamma_{div}L_{div},\vspace{-2.mm}$ (7)
where the first two terms are recognition losses of the downstream gait model, $L_{rec}$ is the mask reconstruction loss, and the last two are the smoothness and diversity losses of the denoising branch. In our implementations, $\gamma_{rec},\gamma_{smo}$ and $\gamma_{div}$ are respectively set to $1.0$, $0.01$, and $5.0$.
@@
Training. All images are processed with Pad-and-Resize to the size of $448\times 224$. We use the SGD optimizer with a weight decay of $0.0005$, a momentum of $0.9$ and an initial learning rate of $0.1$. The learning rate is reduced by a factor of $0.1$ at the $15$k, $25$k, $30$k and $35$k iterations, and the total number of iterations is $40$k. The batch size $(p,k,l)$ is set to $(8,8,30)$, where $p$ is the number of IDs, $k$ the number of sequences per ID, and $l$ the number of frames per sequence. The sampling strategy of frames from sequence follows GaitBase [13]. The data augmentation operation only contains randomly flipping all images in a sequence.
Table 3: Rank-1 accuracy on CCPG: BigGait v.s. recent SoTA models using different inputs. $\text{GaitBase}^{p}$: human parsing inputs. $\text{GaitBase}^{p+s}$: parsing and silhouette inputs.
@@
BigGait CVPR’24 $\mathbf{82.6}$ $\mathbf{85.9}$ $\mathbf{87.1}$ $\mathbf{93.1}$ $\mathbf{87.2}$
Table 4: Rank-1 accuracy on three popular benchmark datasets: BigGait v.s. recent SoTA methods. This is a cross-domain evaluation and comparison, in which all methods are trained on one dataset and tested on the remaining two datasets.
(a) Trained on CCPG
@@
GaitSet [3]
$47.4$
$40.9$
$25.8$
$8.2$
$11.0$
$12.8$
GaitBase [13]
$59.1$
$52.7$
$30.4$
$9.5$
$13.1$
$16.8$
AP3D [18]
$53.7$
$46.2$
$11.9$
$36.2$
$\mathbf{51.6}$
$55.3$
PSTA [46]
$49.7$
$42.3$
$8.8$
$25.7$
$29.4$
$40.6$
BigGait
$\mathbf{77.4}$
$\mathbf{71.5}$
$\mathbf{33.6}$
$\mathbf{43.7}$
$44.8$
$\mathbf{56.4}$
(b) Trained on CASIA-B*
@@
GaitSet
$\mathbf{10.6}$
$16.4$
$17.2$
$24.9$
$7.2$
$12.2$
$12.8$
GaitBase
$\mathbf{10.6}$
$18.1$
$\mathbf{21.4}$
$28.7$
$8.1$
$11.8$
$15.6$
AP3D
$2.1$
$2.9$
$3.9$
$6.1$
$29.3$
$47.4$
$48.3$
PSTA
$1.7$
$1.9$
$3.4$
$5.0$
$19.9$
$37.5$
$34.6$
BigGait
$7.5$
$\mathbf{19.5}$
$14.2$
$\mathbf{43.0}$
$\mathbf{36.9}$
$\mathbf{60.2}$
$\mathbf{64.8}$
(c) Trained on SUSTech1K
@@
GaitSet
$14.0$
$\mathbf{23.7}$
$20.3$
$43.2$
$63.3$
$50.8$
$26.4$
GaitBase
$\mathbf{16.8}$
$21.7$
$\mathbf{26.0}$
$42.7$
$73.1$
$61.2$
$\mathbf{28.2}$
AP3D
$5.5$
$7.9$
$13.9$
$35.1$
$56.7$
$48.1$
$15.3$
PSTA
$3.7$
$5.7$
$9.5$
$26.5$
$31.2$
$27.7$
$10.6$
BigGait
$4.5$
$11.5$
$11.9$
$\mathbf{45.5}$
$\mathbf{91.1}$
$\mathbf{85.8}$
$18.7$
@@ 4.2 Main Results
Note BigGait stands for BigGait-S, as mentioned in Sec. 3.1, where BigGait-L refers to its large version.
@@
Within-domain Evaluation. We compare BigGait with various SoTA methods, such as the skeleton-based [44, 49, 36], silhouette-based [3, 10, 25, 13], and end-to-end gait recognition methods [28], and video-based ReID methods [18, 46, 48], on the challenging CCPG. The Supplementary Material provides more within-domain results on ReID evaluation protocols and on other datasets [47, 39, 54].
@@
As shown in Tab. 3, BigGait achieves significantly better performance than other SoTA methods on CCPG. In the following, we measure BigGait’s effectiveness in filtering out gait-irrelevant noises from two aspects.
@@
First, compared to video-based ReID methods [18, 46, 48], BigGait outperforms them considerably, e.g., $+29.2\%$ for full-changing (CL), $+28.6\%$ for ups-changing (UP), $+17.4\%$ for pants-changing (DN) and $+1.7\%$ for bag-changing (BG) scenarios. Accordingly, we consider that BigGait can efficiently extract robust identity representations from raw RGB videos. Fig. 6 provides intuitive evidence by visualizing the activation difference between BigGait and video-based ReID methods.
@@
Second, compared to silhouette-based gait methods [3, 10, 25, 13], which perform well due to immunity to color/texture noises, BigGait still outperforms them by $+4.0\%$, $+1.1\%$, $+6.4\%$, and $+3.9\%$ in CL, UP, DN, and BG settings. This indicates that BigGait’s immunity to gait-irrelevant noises is comparable to silhouette-based ones.
@@
Based on these analyses, we consider that gait features are the primary characteristics extracted by BigGait, even if gait-irrelevant noises cannot be absolutely eliminated from RGB inputs. We further validate this conclusion through cross-domain and visualization experiments.
@@
Cross-domain Evaluation. Tab. 10 shows cross-domain performance comparisons. BigGait’s cross-domain performance varies depending on the training set used. When trained on CCPG, BigGait exhibits strong adaptability to unseen datasets, i.e., outperforming both video-based ReID methods [18, 46] and silhouette-based methods [13, 3]. When the training set is CASIA-B*, BigGait exihibits less impressive performances in some settings, i.e., performing well in the ups-changing (UP), but poorly in the full-changing (CL) and pants-changing (DN). When crossing domains from SUSTech1K, BigGait struggles in more cross-dressing settings (CL, UP and DN).
@@
We consider that BigGait’s power in filtering gait-irrelevant noises is impacted by the training data biases. Specifically, the probability of sampling cross-dressing data pairs from SUSTech1K is lower than $5.0\%$, likely making BigGait focus on unchanged clothing cues rather than subtle gait patterns. Fortunately, with the enrichment of training data, BigGait can self-correct this mistake. CASIA-B* includes over $35.0\%$ cross-dressing sample pairs, all of which are ups-changing, which improve BigGait’s resistance to ups-changing (UP), as shown in Tab. 10 (b). CCPG has more diverse outfits than others, containing over $95.0\%$ cross-dressing pairs, with $78.5\%$ for ups-changing, $80.0\%$ for pants-changing, and $62.0\%$ for full-changing. As shown in Tab. 10 (a), BigGait can learn robust gait representations by training on CCPG. These analyses indicate that the distribution of training data can significantly influence BigGait’s outcomes, i.e, the more cross-dressing changes given, the more cross-dressing capacity obtained.
Figure 6: Visualization of activation maps. By background removal and texture denoising, BigGait pay more attention on robust gait pattern than video-based ReID methods.
@@
Visualization of Activation Maps. To delve deeper into the gait representation learned by BigGait, Fig. 6 shows visualizations of activation maps from both BigGait and video-based ReID methods [18, 46, 48], driven by the popular Grad-CAM [38] algorithm. These maps are from the first layer in BigGait’s downstream model, second layer in AP3D [18] and PSTA [46], and fourth block in PiT [48]. Four key insights emerge: a) Unlike video-based ReID methods distracted by cluttered backgrounds, BigGait solely attends to the body. b) The video-based ReID methods are sensitive to high-frequency clothing texture, as shown in the red box of Fig. 6. c) $f_{ap}$ and $f_{de}$ target different body regions, with $f_{ap}$ emphasizing limbs and head via parsing-like priors, and $f_{de}$ highlighting human contours by suppressing texture noises on body. Fusing $f_{ap}$ and $f_{de}$ using attention weights to obtain $f_{fusion}$. d) As mentioned before, CASIA-B* only contains ups-changes, and SUSTech1K has few cloth-changes. Thus BigGait incorrectly highlights unchanged pants and clothing when trained on CAISA-B* and SUSTech1K, respectively. More visualizations are provided in Supplementary Material.
@@
4.3 Ablation Study
All experiments are conducted on CCPG. Due to limited GPU resources, we adopt $8$ frames per sequence in the ablation study to train BigGait for more efficient iterations. Despite slight performance drop, BigGait’s performance in Tab. 5 (a) remains comparable with SoTA methods in Tab. 3.
Pad-and-Resize. Compared (a) and (b) in Tab. 5, the Pad-and-Resize strategy is helpful for BigGait by faithfully preserving body aspect ratio.
Various Branch. In Tab. 5 (c), without the GRE module, the performance of BigGait drops to be similar to video-based ReID methods [18] in Tab. 3. This highlights the effectiveness of GRE module. Compared (a) to (d-f) in Tab. 5, each branch contributes substantially. This implies a mutually beneficial relationship among them. In Tab. 5 (g), BigGait’s foreground mask is less discriminable for recognition compared to the silhouette in Tab. 3. We consider that BigGait’s foreground masks is approximate regions rather than precise segmentations. SoTA results in Tab. 5 (a) are achieved even without finely segmented foregrounds.
Denoising Branch. Comparing (a) to (b-d) in Tab. 6, smoothness loss alone leads to a trivial solution, while combining it with diversity constraint improves results. Compared (a) with (e) and (f) in Tab. 6, we explore an appropriate channel number for the denoising branch.
Upstream and Downstream Models. Compared (c) with (d) in Tab. 7, BigGait-L exhibits slightly higher performance than BigGait-S. Compared (c) with (e) in Tab. 7, we verify that the gains primarily come from the knowledge embedded in DINOv2 rather than its model architecture. Compared (c) and (f) in Tab. 7, the performance suffers when fine-tuning the upstream DINOv2. Further, we explore the influence of different upstream/downstream models for BigGait. Compared (c) and (g) in Tab. 7, although the performance drops when changing BigGait’s downstream model, it still surpasses its silhouette-based counterpart in (a). As shown in Tab. 7 (h), BigGait remains effective even when changing BigGait’s upstream model, i.e., replacing DINOv2 [35] with SAM [23]. In most cases, it still outperforms its silhouette-based counterpart in (b). This highlights BigGait’s flexibility for upstream/downstream models.
5 Challenges and Limitations
Challenges for LVMs-based Gait Recognition. Beyond attractive profits brought by BigGait, this paper also reveals two primary challenges for LVMs-based gait recognition: a) Interpretability. While we have introduced several explicit human gait priors as soft constraints, the learned gait representations, in comparison to conventional ones defined by clear and intuitive physical attributes, remain partially understandable. b) Purity. In gait methods that directly utilize RGB videos as input, a recurring challenge involves effectively reducing gait-irrelevant noise within walking sequences. This task becomes even more demanding when attempting to preserve the purity of gait characteristics in LVMs-based gait recognition without explicit supervision.
@@ 5 Challenges and Limitations
[…]
GaitBase $\mathbf{79.0}$ $\mathbf{82.3}$ $\mathbf{86.7}$ $\mathbf{94.5}$
@@
[…]
Limitations of this paper. a) The influence of different upstream LVMs for BigGait is only preliminarily explored. This issue is worth further study. b) The GRE module lacks of spatial-temporal designs. c) This paper focus on high-frequency texture noises instead of low-frequency color noises, since it is the high-frequency that still heavily impacts existing RGB-based methods in the red box of Fig. 6. Further improvements to color noises are expected.